OpenStudy (kellygirl34):

Given the following functions f(x) and g(x), solve f[g(7)] and select the correct answer below: (1 point) f(x) = 4x + 21 g(x) = 2x + 2

OpenStudy (kellygirl34):

the options are 16 49 85 784

OpenStudy (irishboy123):

\[f(x) = 4x + 21\] \[g(x) = 2x + 2\] \[f(g(x)) = f(2x + 2) = 4(2x + 2) + 21\] that's more important than knowing the answer but you should get it from there.....